### v2

# A new approach to delegate signing rights to proxy signers using isogeny-based cryptography

Abstract
E-governance is a two-way protocol through which one can use government services, share data and request information. It refers to the use of communication and information technologies to provide government services to public in an efficient and fast manner. In addition, any document submitted to the e-Government system must be authenticated by a government officer using a digital signature scheme. In the context of digital signatures, the proxy signature is an important cryptographic primitive that allows the original signer to delegate signing authority to another signer (proxy signer). The proxy signature has a number of important applications in the e-government system. There are now a large amount of proxy signature schemes. The security of most of them relies on the following hard problems: the discrete logarithm problem and the factorization of integers problem. However, a large-scale quantum computers can solve them in polynomial time due to Shor’s algorithm. As a consequence, there is a need for a quantum computer-resistant proxy signature to secure e-governance system from quantum adversaries. In this work, we propose the first post-quantum isogeny based proxy signature scheme CSI-PS (commutative supersingular isogeny proxy signature). Our construction is proven to be uf-cma secure under the hardness of the group action inverse problem (GAIP) based on isogeny.
keywords:
E-governance, isogeny-based public-key cryptography; post-quantum cryptography; proxy signature; elliptic curve cryptography.
1 Introduction
Electronic governance often abbreviated as e-governance is an application of information and communication technologies to facilitate the direct interaction between government and other entities like citizens, businesses and employees. In other words, e-governance aims to use technology to achieve the objective of governance. The basic goals of e-governance are to increase individual’s access to information, connect with business and industry, and make government more accountable, transparent, and effective. Essentially, it performs government functions through information, communication, and technology. In e-governance, there are four types of government interactions: government-to-citizen (G2C), government-to-government (G2G), government-to-business (G2B), and government-to-employee (G2E). In e-governance initiatives worldwide, various security measures are implemented to protect the data of residents. Each department’s representatives are provided with digital signature certificates, allowing them to digitally sign documents for authentication purposes. However, a dilemma arises as to how the government representative can sign a paper when he is on leave. That is where proxy signatures come in. A Proxy signature is a type of digital signature, where a participant (original signer) makes some representative (proxy signer) able to sign a message on behalf of the original signer. Given the context of the situation, the proxy signature can be employed by the government official to solve the problem. By making use of such a scheme, he can transfer the signing authority to a proxy signer who can sign on behalf of the government official for a limited time.We demonstrate the above in Figure 1.
Figure 1: Proxy signature in e-governance
The proxy signature idea was developed by Mambo et al. [1] in 1996. A proxy signature should attain the below-mentioned security features [2]:
  1. 1.
    Unforgeability: No one but the proxy signer, not even the original signer, should be able to generate a correct proxy signature.
  2. 2.
    Identifiability: The proxy signer should be identified after observing the proxy signature by anyone having the authority to access the signature.
  3. 3.
    Undeniability: After a proxy signer generates a valid proxy signature for the original signer, they cannot later deny having created it.
  4. 4.
    Verifiability: The verifier can be convinced of the original singer’s agreement from the proxy signature.
  5. 5.
    Distinguishability: In polynomial time, the original signer’s valid signature can be distinguished from a valid proxy signature.
  6. 6.
    Secrecy: Any information about the secret key of the original signer cannot be determined at any stage during proxy signature.
  7. 7.
    Prevention of misuse: Proxy signer will not be able to sign a message with some proxy key that is not defined in its warrant. Otherwise, it will be identified from its warrant.
  8. 8.
    Revocability: The proxy signer’s signing authority will be disabled automatically on completion of the validity of the period. Then any proxy signature produced by the proxy signer will not be valid.
In addition to applications in e-governance, proxy signature find applications in other real-life situations like global distribution networks [3, 4], mobile agent applications [5, 6, 7], etc. The majority of the currently used proxy signature schemes [8, 9, 10, 11, 12, 13] are based on hard problems like the discrete logarithm problem and the integer factorization problem. However, they are not protected against attacks by quantum computers since Shor’s algorithm [14] can efficiently solve them. Consequently, post-quantum cryptography (PQC) is considered a promising alternative believed to offer security against quantum adversaries. There are several candidates of PQC, such as Lattice-based cryptography [15], Multivariate cryptography [16], Isogeny-based cryptography [17], Hash-based cryptography [18], Code-based cryptography [19]. Isogeny-based cryptography (IBC) is a relatively new area within post-quantum cryptography (PQC) that offers small public key sizes compared to other PQC candidates. The concept of IBC was first introduced by Couveignes [20] in 1997. That work was based on ordinary elliptic curves for which the endomorphism ring is an order of some imaginary quadratic field (IQF). In 2006, Rostovtsev and Stolbunov [21] rediscovered the idea and proposed a Diffie-Hellman [22] like key exchange (KE) protocol. In 2018, Castryck et al.  [23] developed an efficient key agreement protocol ${\sf Commutative~{}supersingular~{}Isogeny~{}Diffie-Hellman~{}(CSIDH)}$ which uses supersingular elliptic curves defined over a prime finite field. However, instead of the full endomorphism ring, they worked with the $\mathbb{F}_{p}$-rational endomorphism ring, which is an order of IQF and hence commutative. Throughout the literature, the emphasis is on isogeny-based protocols relying on class group action [17]. Research shows that there are many advantages [24, 25] in isogeny-based cryptography, such as the size of the signatures and public keys, which are almost the same as existing practical schemes. Another major advantage is its sufficiently low communication cost. In the literature, there are some efficient isogeny based digital signatures like ‘SeaSign’ [26], ‘CSI-Fish’[27], ‘SQISign’ [28] which are based on the ‘Fiat-Shamir’ technique. Additionally, there are isogeny-based identity based signature scheme (IBS) [29], Group signature [30, 31] and Ring signature [30] and Signcryption [32, 33]. However, there is a lack of proxy signatures based on IBC. Consequently, designing a secure and efficient isogeny-based proxy signature is an interesting open challenge.
1.1 Our contribution
The major contributions of the paper are outlined below.
  • To the extent of our knowledge, there has been no construction of proxy signature in the context of isogeny-based cryptography. In this paper, we addressed this open challenge by designing the first isogeny-based post-quantum secure proxy signature CSI-PS (Commutative Supersingular Isogeny Proxy Signature).
  • The security of our proposed design CSI-PS is based on hardness assumption of the group action inverse problem ($\sf{GAIP}$). In particular, we show that CSI-PS achieves uf-cma security under the random oracle model. Moreover, CSI-PS also meets the other security requirements, such as identifiability, undeniability, and verifiability.
  • We observe that the public keys of original signer and proxy signer are of sizes $L_{0}\left\lceil{log(p)}\right\rceil$ bits, original signer’s and proxy signer’s secret key sizes are $L_{0}\left\lceil{nlog(2I_{0}+1)}\right\rceil$ bits each, sizes of proxy share and proxy signature become $M_{1}M_{2}\left\lceil{log(L_{1}+1)}\right\rceil$ + $M_{1}M_{2}\left\lceil{nlog(2I_{1}+1)}\right\rceil$ bits each.
  • To determine the most optimal size for the proxy signature, we perform calculations for different parameter choices. We conclude that our proxy signature achieves a size of 3052B with parameters set to $(16,8,8,2,65535,255)$. A concrete instantiation of CSI-PS for the parameters set $(16,8,8,2,65535,255)$ is provided in Section 6. We demonstrate how an original signer can use our proposed CSI-PS design to create a proxy signer with signing authority, and how the proxy signer can then generate a valid proxy signature on a document.
  • The only cost intensive operation in our design is class group computation. We computed the cost of class group computation in terms of field multiplications. In addition, we calculated the exact number of field multiplications required for public key generation, proxy share generation, and proxy signature. Moreover, we also estimated the time complexity of CSI-PS. The results are summarized in Table 3.
1.2 Organization of the paper
The structure of the paper is outlined as follows: Section 2 covers the preliminaries to aid readers in understanding our scheme. Section 3 introduces our scheme, CSI-PS. Security analysis is discussed in Section 4, while communication complexities are examined in Section 5. A concrete instantiation illustrating the CSI-PS flow is provided in Section 6. Finally, we conclude our discussion in Section 7.
2 Preliminaries
Notations. Given a finite set $X$, $|X|$ stands for its cardinality. $\mathbb{F}_{q}$ denotes the finite field of order $q$, where $q$ is a power of some prime $p$. Closure of a field $\mathbb{F}_{q}$ is denoted by $\overline{\mathbb{F}}_{q}$. $[a,b]$ denotes the set of integers $z$ such that $a\leq z\leq b$.
Definition 1.
Negligible function:
A function $\epsilon:\mathbb{N}\longrightarrow\mathbb{R}$ is called a negligible function, if for every polynomial poly(x) in $x$, there exists a positive integer $N$ such that $\epsilon(x)<\frac{1}{poly(x)}$ $\forall$ $x>N$
2.1 Elliptic curves and isogenies [17]
An elliptic curve over a field $\mathbb{F}_{q}$ is a smooth projective curve of genus $1$ having at least one $\mathbb{F}_{q}$-rational point. For char$(\mathbb{F}_{q})\not=2,3$, the short Weierstrass affine equation for an elliptic curve over $\mathbb{F}_{q}$ is $y^{2}=x^{3}+cx+d$, where the point at infinity is $(0:1:0)$, and $c,d\in\mathbb{F}_{q}$ with $04c^{3}+27d^{2}\not=$ 0. Throughout the paper, elliptic curves are assumed to be defined over $\mathbb{F}_{p}$. According to Hasse’s theorem [34], the cardinality $|E(\mathbb{F}_{p})|$ of $E(\mathbb{F}_{p})$ satisfies $|E(\mathbb{F}_{p})|=p+1-t$ with $|t|\leq 2\sqrt{p}$.
An isogeny $\chi:E_{1}\rightarrow E_{2}$ is an onto morphism inducing a group homomorphism from $E_{1}(\overline{\mathbb{F}}_{q})$ to $E_{2}(\overline{\mathbb{F}_{q}})$. Here $E_{1}$ and $E_{2}$ are called isogenous. Indeed, $E_{1}$ and $E_{2}$ having the same number of points assure an isogeny between them. An isogeny sends the identity element $\Theta$ of $E_{1}$ to identity element $\Theta^{\prime}$ of $E_{2}$. An endomorphism is a morphism from $E$ to itself that fixes the distinguished point $\Theta$. The set of all endomorphisms on $E$ is denoted by $\text{End}(E)$. One should say that the set of all endomorphisms is a ring. The Frobenius endomorphism is an endomorphism on $E$ over $\mathbb{F}_{q}$ and it is defined as $\pi_{E}:(x:y:z)\longmapsto(x^{q}:y^{q}:z^{q})$. Another example of an endomorphism is the multiplication-by-$n$ map, defined as $[n]:P\longmapsto[n]P$. Its kernel is called the $n$-th torsion subgroup $E[n]$.
The standard form of isogeny $\chi$ over $\mathbb{F}_{q}$ is given by $\chi(x,y)=\left(\frac{\alpha(x)}{\beta(x)},\frac{\gamma(x)}{\delta(x)}y\right)$, where $\alpha,\beta,\gamma,\delta\in\mathbb{F}_{q}[x]$, $\alpha$ is co-prime to $\beta$ and $\gamma$ is co-prime to $\delta$ in $\mathbb{F}_{q}[x]$. The degree of $\chi$ is denoted by $deg(\chi)$ and defined by $deg(\chi)=max\{deg(\alpha),deg(\beta)\}$. If $\left(\frac{\alpha(x)}{\beta(x)}\right)^{\prime}\neq 0$ then $\chi$ is called separable; otherwise, it is called inseparable. Note that the Frobenius endomorphism over $\mathbb{F}_{p}$ is an inseparable isogeny. The endomorphism $[m]$ for an elliptic curve $E/\mathbb{F}_{p}$ is separable if and only if $m$ and $p$ are co-prime. The $j$-invariant of $y^{2}=x^{3}+cx+d$ is defined by $j(E)=1728\frac{4c^{3}}{4c^{3}+27d^{2}}$. Two elliptic curves over $\overline{\mathbb{F}_{q}}$ with the same $j$-invariant are called isomorphic. The torsion subgroup $E[p]$ is isomorphic to either $\mathbb{Z}/p\mathbb{Z}$ or 0. E is called ordinary in the first case, while $E$ is called a supersingular elliptic curve in the second case. If $E$ is supersingular over a finite field of characteristic $p$, then $j(E)$ lies in $\mathbb{F}_{p^{2}}$.
2.2 Endomorphism algebra and ideal class group [17]
The endomorphism ring $\text{End}(E)$ is not just a $\mathbb{Z}$-module but also a $\mathbb{Z}$-algebra. The endomorphism algebra of $E$ is denoted by $\text{End}^{0}(E):=\text{End}(E)\otimes_{\mathbb{Z}}\mathbb{Q}$, making $\text{End}^{0}(E)$ a $\mathbb{Q}$-algebra. Since $\text{End}(E)$ is a free $\mathbb{Z}$-module of rank $r$, note the following cases: if $r=1$, then $\text{End}^{0}(E)$ is isomorphic to $\mathbb{Q}$; if $r=2$, then $\text{End}^{0}(E)$ is isomorphic to an imaginary quadratic field; and if $r=4$, then $\text{End}^{0}(E)$ is isomorphic to a quaternion algebra. If $E$ is defined over a finite field, the case $r=1$ cannot occur. An order $\vartheta$ of a $\mathbb{Q}$-algebra $K$ is a subring of $K$ that has the same rank as $K$. For any number field $K$, the ring of integers $\vartheta_{K}$ is the unique maximal order. In imaginary quadratic number fields, orders are defined by $\vartheta=\mathbb{Z}+f\vartheta_{K}$, where $f$ is an integer called the conductor of $\vartheta$.
The Frobenius endomorphism $\pi$ for an elliptic curve $E/\mathbb{F}_{q}$ satisfies the equation $\pi^{2}-t\pi+q=0$ with $|t|\leq 2\sqrt{q}$. Here $t$ is called trace of $\pi$ and $\Delta_{\pi}=t^{2}-4q$ is discriminant of $\pi$. Therefore, the discriminant $\Delta_{\pi}=t^{2}-4q$ of $\pi$ is negative and we have $\pi\in\mathbb{Q}(\sqrt{\Delta_{\pi}})$ which is an imaginary quadratic field. Let $\vartheta$ be the order of a number field $K$. With the help of an $\vartheta$-ideal $L$, we can define a fractional $\vartheta$- ideal as $L^{\prime}=\omega L=\{\omega\alpha:\alpha\in L\}$ for some $\omega\in K^{*}$. A fractioal $\vartheta$-ideal $L^{\prime}$ is called invertible if there exists a fractional $\vartheta$-ideal $L^{\prime\prime}$ such that $L^{\prime}L^{\prime\prime}=\vartheta$. Collection of invertible fractional $\vartheta$-ideals forms a group with respect to multiplication of fractional $\vartheta$-ideals which is defined as $L^{\prime}N^{\prime}=\omega\omega^{\prime}LN$ for two fractional $\vartheta$-ideals $L^{\prime}=\omega L$ and $N^{\prime}=\omega^{\prime}N$. Let $I$ represent the group of invertible fractional $\vartheta$-ideals and $P$ denotes the subgroup of $I$ consisting of principal fractional $\vartheta$-ideals. The ideal class group is defined by the quotient group $cl(\vartheta)=\frac{I}{P}$. The $\mathbb{F}_{p}$-rational endomorphism ring $\text{End}_{p}(E)$ for a supersingular elliptic curve $E$ over $\mathbb{F}_{p}$ is an order in an imaginary quadratic field. Suppose $E_{p}$ is the set of isomorphism classes of elliptic curves over $\mathbb{F}_{p}$ whose $\mathbb{F}_{p}$-rational endomorphism ring is isomorphic to $\vartheta$. Then, we have a group action of $Cl(\vartheta)$ on $E_{p}$ which is free and transitive.
Theorem 1.
Let $E$ be a supersingular elliptic curve defined over $\mathbb{F}_{p}$ with the Frobenius endomorphism $\pi$, where $p\equiv 3\ (\textrm{mod}\ 4)$ and $p>3$. The necessary and sufficient condition for $\text{End}_{p}(E)=\mathbb{Z}[\pi]$ is that there exists a unique $e\in\mathbb{F}_{p}$ so that $E$ is $\mathbb{F}_{p}$ isomorphic to a Montgomery curve: $y^{2}=x^{3}+ex^{2}+x$, where $\mathbb{Z}[\pi]$ is a subring of $\mathbb{Q}(\sqrt{-p})$.
Algorithm 1 The class-group action on a Montgomery curve
1:A Montgomery curve $E:y^{2}=x^{3}+ex^{2}+x$, $e\in\mathbb{F}_{p}$, $p=4l_{1}\cdots l_{n}-1$ and a set of integers $(a_{1},\cdots,a_{n})$.
2:while some $a_{i}\neq 0$ do
3:     Sample $x\in\mathbb{F}_{p}$.
4:     Set $w\longleftarrow 1$ if $x^{3}+ex^{2}+x$ is a square in $\mathbb{F}_{p}$, else set $w\longleftarrow-1$.
5:     Suppose $W=\{i:a_{i}\neq 0,sign(a_{i})=w\}$.
6:     if $W=\emptyset$ then
7:         Start from step 2.
8:     end if
9:     Set $f\longleftarrow\prod_{i\in W}l_{i}$.
10:     Evaluate $Q\longleftarrow[(p+1)/f]P$, where $P$ is a point of order $l_{i}$ in E.
11:     for each $i\in W$ do
12:         Calculate $R\longleftarrow[f/l_{i}]Q$.
13:         if $R=\infty$ then
14:              skip
15:         end if
16:         Compute the isogeny $\phi:E\longrightarrow E^{\prime}$ with $E^{\prime}:y^{2}=x^{3}+Bx^{2}+x$ and $ker(\phi)=R$.
17:         Set $Q\longleftarrow\phi(Q)$, $f\longleftarrow f/l_{i}$, $a_{i}=a_{i}-w$
18:     end for
19:end while
20:$B$.
2.3 Commutative supersingular isogeny Diffie-Hellman (CSIDH) [23]
If $E$ is a supersingular curve over $\mathbb{F}_{p}$ then its ring of $\mathbb{F}_{p}$-rational endomorphism is an imaginary quadratic order $\vartheta\subseteq\mathbb{Q}(\sqrt{-p})$ which is denoted by $\text{End}_{p}(E)$. In CSIDH [23], we fix $E_{0}:y^{2}=x^{3}+x$ over $\mathbb{F}_{p}$ for the prime $p=4l_{1}\cdots l_{n}-1$, where $l_{1},\cdots,l_{n}$ are small distinct odd primes. As $p\equiv(-1)(\bmod~{}4)$, we can say that $E_{0}$ is supersingular. Hence $|E_{0}(\mathbb{F}_{p})|=p+1\equiv 0(\bmod~{}l_{i})$ and we can decompose $l_{i}\vartheta=J_{i}\overline{J_{i}}$, where $J_{i}=(l_{i},\pi-1)$ and $\overline{J_{i}}=(l_{i},\pi+1)$. Now we have to assume that $J_{i}$’s are evenly distributed in the class group and we expect that an ideal is of the form $J_{1}^{a_{1}}\cdots J_{n}^{a_{n}}$ with $(a_{1},\cdots a_{n})$ randomly chosen from $[-s,s]^{n}$ for some natural number $s$. Thereby, we may represent the ideal $J_{1}^{a_{1}}\cdots J_{n}^{a_{n}}$ as $(a_{1},\cdots,a_{n})$. It has been observed that it is sufficient to choose $m$ such that $2s+1\geq\sqrt[n]{|cl(\vartheta)}$. As the trace of Frobenius endomorphism is zero for supersingular elliptic curve over $\mathbb{F}_{p}$, its characteristic equation is $\pi^{2}+p=0$ yielding $\pi=\sqrt{-p}$. Thus, $\text{End}_{p}(E_{0})=\mathbb{Z}[\sqrt{-p}]=\mathbb{Z}[{\pi}]$. In other words, $E_{0}$ is $\mathbb{F}_{p}$ rational isomorphic to a Montgomery curve $E_{A}:y^{2}=x^{3}+ex^{2}+x$, which is represented by its coefficient $e$ by theorem 1.
To generate the key, Alice randomly chooses a n-tuple secret key $\textbf{a}=(a_{1},\cdots,a_{n})$ randomly form $[-s,s]^{n}$ and computes $[\textbf{a}]=[J_{1}^{a_{1}}\cdots J_{n}^{a_{n}}]\in\mathbb{Z}[\pi]$. She then computes the group action $[\textbf{a}]E_{0}$ using Algorithm 1, which is isomorphic to some Montgomery curve $E_{A}$ with a coefficient $f$ (public key) by Theorem 1. Here, Alice’s secret key - public key pair is $([\textbf{a}],f)$. Similarly Bob chooses his key pairs as $([\textbf{b}],g)$. Next, Alice evaluates $[\textbf{a}]E_{B}=[\textbf{a}]([\textbf{b}]E_{0})$ and Bob computes $[\textbf{b}]E_{A}=[\textbf{b}]([\textbf{a}]E_{0})$. At the end of the protocol, they will share the same key $[\textbf{a}][\textbf{b}]E_{0}=[\textbf{b}][\textbf{a}]E_{0}$ due to the commutative property of imaginary quadratic field.
2.4 Rejection sampling [29]
In order to prevent any leakage of information of the secret key, we can apply rejection sampling. Let there be two parties P and Q. P selects a secret vector $\textbf{a}=(a_{1},\cdots,a_{n})$in $[-I,I]^{n}$, where $I\in\mathbb{N}$ so that $\prod_{t=1}^{n}J_{t}^{a_{t}}$ covers almost all the ideal classes in order to make the output distribution uniform. In addition, P chooses another random vector $\textbf{a}^{\prime}$in [-($\delta$ + 1)$I$ , ($\delta$ + 1)$I$]n where $\delta\in\mathbb{N}$. After that, Q randomly chooses a bit $b$ and sends it to P. In the following, P evaluates ${\gamma}=\textbf{a}^{\prime}-\textbf{a}$ if ${b}=1$ and otherwise, $\gamma=\textbf{a}^{\prime}$. As a consequence, we can filter $\gamma$ as $|\gamma|\leq\delta I$ to prevent any leakage of the secret key. Here the output vector $\gamma$ is uniformly distributed and as a result it is independent of a.
2.5 Hardness assumption
Definition 2.
Group Action Inverse Problem ($\sf{GAIP}$[26]:
[…]
2.6 SeaSign [26]
The signature scheme SeaSign [26] consists of three algorithms ${\sf Key~{}generation}$, ${\sf Signature}$ and ${\sf Verification}$ which are discussed below.
$({PK},{SK})\leftarrow{\sf Key~{}generation}(1^{\lambda}):$ Consider a supersingular elliptic curve $E_{0}:y^{2}=x^{3}+x$ over $\mathbb{F}_{p}$ where $p$ is a prime. On input of a security parameter $\lambda$, the signer randomly chooses a=$(a_{1},\cdots,a_{n})$in $[-I_{0},I_{0}]^{n}$ and evaluates $E_{1}=[\textbf{a}]E_{0}$. The public key-secret key pair of the signer is $({PK,SK})=(E_{1},\textbf{a})$.
$(\sigma)\leftarrow{\sf Signature}(m,{SK},{PK}):$ On input of a message $m$, the signer performs the following steps:
  1. 1.
    chooses $\textbf{x}^{i}\in_{R}[-(I_{0}+I_{1}),(I_{0}+I_{1})]^{n}$ and evaluates $X_{i}=[\textbf{x}^{i}]E_{0}$ for $i=1,2,\ldots,t$,
  2. 2.
    calculates $H(X_{1},\cdots,X_{t},m)=b_{1}\parallel\cdots\parallel b_{t}$, where $H:\{0,1\}^{*}\rightarrow\{0,1\}^{t}$ is a cryptographically secure collision resistant hash function,
  3. 3.
    sets $\textbf{z}^{k}=\textbf{x}^{k}$ if $b_{k}=0$ and $\textbf{z}^{k}=\textbf{x}^{k}-\textbf{a}$ otherwise by ensuring the fact that $\textbf{z}^{k}\in[-I_{1},I_{1}]^{n}$,
  4. 4.
    outputs the signature as $\sigma=(z_{1},z_{2},\cdots,z_{t},b_{1},b_{2},\cdots,b_{t})$.
$(1~{}\mbox{or}~{}0)\leftarrow{\sf Verification}(m,\sigma,{PK}):$ On receiving a signature $\sigma$ from the sender, a verifier executes the following operations:
  1. 1.
    computes
    $X^{\prime}_{k}=\left.\begin{cases}[z_{k}]\mathcal{E},&\text{if }b_{k}=0\\ [z_{k}]\mathcal{E}_{1},&\text{if }b_{k}=1\par\end{cases}\right\},$
  2. 2.
    evaluates $H(X^{\prime}_{1},\cdots,X^{\prime}_{t},m)=b^{\prime}_{1}\parallel\cdots\parallel b^{\prime}_{t}$ ,
  3. 3.
    outputs 1 if the equality $b_{1}\parallel\cdots\parallel b_{t}=b^{\prime}_{1}\parallel\cdots\parallel b^{\prime}_{t}$ holds, else outputs 0.
Theorem 2.
[26] The above signature scheme is unforgeable under chosen-message attack (uf-cma) relying on the hard problem $\sf{GAIP}$ in the random oracle model.
2.7 Syntax of a proxy signature scheme [9]
In general, a proxy signature involves two entities A (original signer) and B (proxy signer) with the following five algorithms:
(pp) $\leftarrow$ ${\sf Setup}$ $(1^{\lambda})$:- On input the security parameter $\lambda\in\mathbb{N}$, this algorithm outputs the collection of public parameters pp which are published in a public platform like a bulletin board.
${((PK_{A},SK_{A}),(PK_{B},SK_{B}))}$leftarrow ${\sf Key~{}generation}$ (pp):- In this algorithm, A and B generate their key pairs $(PK_{A},SK_{A})$ and $(PK_{B},SK_{B})$ respectively.
(${z_{B}}$) $\leftarrow$ ${\sf Proxy~{}share~{}generation}$ ${(SK_{A},d_{B})}$:- Let $d_{B}$ be the attribute of B consisting of the necessary information like DOB, id card number, etc. Then A generates a proxy share $z_{B}$ with the help of $SK_{A}$ in favor of B and publishes $(d_{B},z_{B})$ as a warrant in the bulletin board.
(0 or 1) $\leftarrow$ ${\sf Proxy~{}share~{}verification}$ $(PK_{A},d_{B},z_{B})$:- B runs this algorithm to check the validity of the warrant $(d_{B},z_{B})$. The algorithm outputs 1 if the warrant is valid; otherwise, it outputs 0.
($\sigma$) $\leftarrow$ ${\sf Proxy~{}signature}$ $(m,SK_{B},d_{B},z_{B})$:- The proxy signer B holding a valid warrant, outputs a signature $\sigma$ on some message using its secret key $SK_{B}$.
(0 or 1) $\leftarrow$ ${\sf Proxy~{}signature~{}verification}$ $(m,\sigma,PK_{A},PK_{B},d_{B},z_{B})$:- Given a message-signature pair $(m,\sigma)$ along with a warrant $(d_{B},z_{B})$ of the proxy signer B, the verifier checks the validity of the warrant $(d_{B},z_{B})$. If the warrant is valid, the verifier checks the correctness of $(m,\sigma)$ using $PK_{B}$. It then outputs a deterministic value of 1 if the verification is done correctly; else, it outputs 0.
2.8 Existential unforgeability under chosen-message attack (uf-cma)
Let us consider a proxy signature scheme consisting of the algorithms: Setup, Key generation, Proxy share generation, Proxy share verification, Proxy signature, Proxy signature verification. The uf-cma security for the proxy signature scheme is defined by a game between an adversary ($\mathcal{A}$) and a challenger ($\mathcal{C}$) for a particular proxy signer with attribute $d$. Once $d$ is chosen, it remains unchanged throughout the game. In this game $\mathcal{C}$ generates the key pair for original signer and the proxy signer. In the following, $\mathcal{C}$ responds to the queries made by $\mathcal{A}$ over some messages polynomial number of times with $d^{\prime}$. Finally, the $\mathcal{A}$ has to output a valid message-signature pair $(m^{\star},\sigma^{\star})$ to win the game, where the message $m^{\star}$ should not be queried before. We denote this experiment by $Exp^{uf-cma}_{proxy(1^{\lambda})}$ which is described below in detail.
Setup: $\mathcal{C}$ runs the Key generation algorithm on input pp to generate a new public key-secret key pairs $(PK_{O},SK_{O})$ for original signer and $(PK_{P},SK_{P})$ for proxy signer on input of security parameter $\lambda$.
Sign-query: In this phase, $\mathcal{A}$ queries to $\mathcal{C}$ of proxy signature for a message m with the attribute $d$ of the proxy signer. $\mathcal{C}$ first recognizes the warrant $(d,z)$ for $d$ from the bulletin board and runs the algorithm Proxy signature to generate a proxy signature $\sigma$ which is sent to $\mathcal{A}$. $\mathcal{A}$ can query polynomial number of times.
Forgery: $\mathcal{A}$ outputs a forge signature $\sigma^{\star}$ for a message $m^{\star}$ which has not been previously queried. $\mathcal{A}$ succeeds if it can impersonate the proxy signer i.e., Proxy signature verification outputs 1 on input $(m^{\star},\sigma^{\star},PK_{O},PK_{P},d,z)$.
The advantage of $\mathcal{A}$ is defined as $Adv_{\mathcal{A}}^{Exp^{uf-cma}_{proxy(1^{\lambda})}}$ = Prob[$Exp^{uf-cma}_{proxy(1^{\lambda})}$ = 1] = Prob[$\mathcal{A}$ wins the game].
Definition 3.
A prosy signature scheme is said to be uf-cma secure if for any probabilistic polynomial time adversary $\mathcal{A}$,
$Adv_{\mathcal{A}}^{Exp^{uf-cma}_{proxy(1^{\lambda})}}\leq\epsilon(\lambda)$
where $\epsilon(\lambda)$ is a negligible function of $\lambda$.
3 Proposed isogeny based proxy signature
In this section, we describe our proposed scheme CSI-PS, where the original signer (A) delegates its signing authority to the proxy signer (B). Our approach consists of six algorithms: Setup, key generation, Proxy share generation, Proxy share verification, Proxy signature, Proxy signature verification. During the ${\sf Setup}$ phase, the original signer generates the public parameters. Key generation algorithm enables A and B to generate their public key-secret key pairs. During the Proxy share generation phase, A will generate a proxy share $z_{B}$ for B by using its secret key and B’s attribute $d_{B}$ consisting of identity and other necessary information, like DOB, identity card number, etc. of B. On receiving a proxy share $z_{B}$ from A, the proxy signer B checks the validity of the warrant $(d_{B},z_{B})$ by implementing the algorithm Proxy share verification. During the Proxy signature algorithm, B generates a signature on a message $m$ with its secret key and proxy share $z_{B}$. The verifier verifies the correctness of the warrant and validity of the message-signature pair during the Proxy signature verification phase. A comprehensive description is provided below:
[…]
${((PK_{A},SK_{A}),(PK_{B},SK_{B}))}$leftarrow ${\sf Key~{}generation}$ (pp):- A does the following operations for $i=1,2,\cdots,L_{0}$:
  1. 1.
    Randomly chooses $\textbf{a}^{(i)}$=$(a_{1}^{(i)},\cdots,a_{n}^{(i)})$in $[-I_{0},I_{0}]^{n}$.
  2. 2.
    Computes $E_{i}=[\textbf{a}^{(i)}]E_{0}.$
  3. 3.
    Publishes the public key $PK_{A}={(E_{i})}^{L_{0}}_{i=1}$ and keeps the secret key $SK_{A}={(\textbf{a}^{(i)})}^{L_{0}}_{i=1}.$
On the other hand, B executes the following for $i=1,2,\cdots,L_{0}$:
  1. 1.
    Randomly selects $\textbf{b}^{(i)}$=$(b_{1}^{(i)},\cdots,b_{n}^{(i)})$in $[-I_{0},I_{0}]^{n}$.
  2. 2.
    Evaluates $\hat{E_{i}}=[\textbf{b}^{(i)}]E_{0}$.
  3. 3.
    Sets the public key as $PK_{B}={(\hat{E_{i}})}^{L_{0}}_{i=1}$ and the secret key as $SK_{B}={(\textbf{b}^{(i)})}^{L_{0}}_{i=1}$.
Algorithm 2 Key generation (original signer)
1:pp
2:for each $i\in 1,\cdots,L_{0}$ do
3:     $\textbf{a}^{(i)}$=$(a_{1}^{(i)},\cdots,a_{n}^{(i)})$ $\in_{R}$ $[-I_{0},I_{0}]^{n}$.
4:     Evaluate $E_{i}=[\textbf{a}^{(i)}]E_{0}$.
5:end for
6:$PK_{A}={(E_{i})}^{L_{0}}_{i=1}$, $SK_{A}={(\textbf{a}^{(i)})}^{L_{0}}_{i=1}$.
Algorithm 3 Key generation (proxy signer)
1:pp
2:for each $i\in 1,2,\cdots,L_{0}$ do
3:     $\textbf{b}^{(i)}$=$(b_{1}^{(i)},\cdots,b_{n}^{(i)})$ $\in_{R}$ $[-I_{0},I_{0}]^{n}$.
4:     Evaluate $\hat{E_{i}}=[\textbf{b}^{(i)}]E_{0}$.
5:end for
[…]
(${z_{B}}$) $\leftarrow$ ${\sf Proxy~{}share~{}generation}$ ${(SK_{A},d_{B})}$:- The original signer A executes the following steps to generate a proxy share $z_{B}$ on input $SK_{A}$ and B’s attribute $d_{B}$. To do that, A performs the following tasks.
  1. 1.
    Computes $h=H_{1}(d_{B},pp)$, where $H_{1}:\{{0,1}\}^{*}\rightarrow(\mathbb{Z}_{{L_{0}}+1})^{M_{1}}$ is a cryptographically secure collision resistant hash function and parses $h$ as integer $\{h_{i}\}_{i=1}^{M_{1}}$.
  2. 2.
    For $i=1,\cdots,M_{1}$ and $j=1,\cdots,M_{2}$ chooses $\textbf{x}^{(i,j)}\in_{R}[-(I_{0}+I_{1}),(I_{0}+I_{1})]^{n}$ and evaluates, $X_{ij}=[\textbf{x}^{(i,j)}]E_{h_{i}}$.
  3. 3.
    Determines $C=H_{2}(d_{B},pp,\{{X}_{ij}\}_{i=1j=1}^{M_{1}M_{2}})$ for cryptographically secure collision resistant hash function $H_{2}:\{{0,1}\}^{*}\rightarrow(\mathbb{Z}_{{L_{1}}+1})^{M_{1}M_{2}}$ and parses it to integers $\{C_{(i,j)}\}_{i=1j=1}^{M_{1}M_{2}}$.
  4. 4.
    For $i=1,\cdots,M_{1}$ and $j=1,\cdots,M_{2}$, sets
    $\textbf{y}^{(i,j)}=\begin{cases}\textbf{x}^{(i,j)},&\text{if $C_{(i,j)}=0$}\\ \textbf{x}^{(i,j)}+\textbf{a}^{(h_{i})},&\text{otherwise}\end{cases}$
  5. 5.
    If $\textbf{y}^{(i,j)}\in[-I_{1},I_{1}]^{n}$ then performs the next step; otherwise, starts from step 2.
  6. 6.
    Outputs the proxy share as $z_{B}=(C,\{\mathbf{y}^{(i,j)}\}^{M_{1},M_{2}}_{i=1j=1})$ which is delivered to B and publishes $(d_{B},z_{B})$ as warrant on the bulletin board.
Algorithm 4 Proxy share generation
1:${(SK_{A},d_{B})}$, $d_{B}$ is some attributes of $B$.
2:Computes $h=H_{1}(d_{B},pp)$. $H_{1}:\{{0,1}\}^{*}\rightarrow(\mathbb{Z}_{{L_{0}}+1})^{M_{1}}$ is a hash function.
3:Parses $h$ as integer $\{h_{i}\}_{i=1}^{M_{1}}$.
4:for each $i=1,\cdots,M_{1}$ and $j=1,\cdots,M_{2}$ do
5:     Chooses $\textbf{x}^{(i,j)}\in_{R}[-(I_{0}+I_{1}),(I_{0}+I_{1})]^{n}$.
6:     Evaluates $X_{ij}=[\textbf{x}^{(i,j)}]E_{h_{i}}$.
7:end for
8:$C=H_{2}(d_{B},pp,\{{X}_{ij}\}_{i=1j=1}^{M_{1}M_{2}})$ for a cryptographically secure collision resistant hash function $H_{2}:\{{0,1}\}^{*}\rightarrow(\mathbb{Z}_{{L_{1}}+1})^{M_{1}M_{2}}$.
9:Parses $C$ to a integer set $\{C_{(i,j)}\}_{i=1j=1}^{M_{1}M_{2}}$.
10:for each $i=1,\cdots,M_{1}$ and $j=1,\cdots,M_{2}$ do
11:     if $C_{(i,j)}=0$ then
12:         $\textbf{y}^{(i,j)}=\textbf{x}^{(i,j)}$
13:     else
14:         $\textbf{y}^{(i,j)}=\textbf{x}^{(i,j)}+\textbf{a}^{(h_{i})}$
15:     end if
16:     if $\textbf{y}^{(i,j)}\in[-I_{1},I_{1}]^{n}$ then
17:         Perform the next step.
18:     else
19:         Start from Step 3.
20:     end if
21:end for
22:$z_{B}=(C,\{\mathbf{y}^{(i,j)}\}^{M_{1},M_{2}}_{i=1j=1})$ and publishes $(d_{B},z_{B})$ as the warrant.
(0 or 1) $\leftarrow$ ${\sf Proxy~{}share~{}verification}$ $(PK_{A},d_{B},z_{B})$:- On input $(PK_{A},d_{B},z_{B})$, the proxy signer B performs the following tasks to check whether the warrant $(d_{B},z_{B})$ is valid or not:
  1. 1.
    Determines $h=H_{1}(d_{B},pp)$, and parses $h$ and $C$ into integers as $\{h_{i}\}_{i=1}^{M_{1}}$ with $0\leq h_{i}\leq L_{0}$ and $\{C_{(i,j)}\}_{i=1j=1}^{M_{1}M_{2}}$ with $0\leq C_{(i,j)}\leq L_{1}$ respectively.
  2. 2.
    For $i=1,\cdots,M_{1}$ and $j=1,\cdots,M_{2}$ sets
    $\overline{X}_{ij}=\begin{cases}[\textbf{y}^{(i,j)}]E_{h_{i}},&\text{if $C_{(i,j)}=0$}\\ [\textbf{y}^{(i,j)}]E_{0},&\text{otherwise}\end{cases}$
  3. 3.
    Determines $C^{\prime}=H_{2}(d_{B},pp,\{\overline{X}_{ij}\}_{i=1j=1}^{M_{1}M_{2}})$.
  4. 4.
    If $C=C^{\prime}$, outputs 1 and gets the assurance of signing authority on behalf of the original signer A; otherwise, outputs 0.
Algorithm 5 Proxy share verification
1:$(PK_{A},d_{B},z_{B})$.
2:Evaluate $h=H_{1}(d_{B},pp)$.
3:Parses $h$ as integers $\{h_{i}\}_{i=1}^{M_{1}}$, $0\leq h_{i}\leq L_{0}$.
4:Parses $C$ as integers $\{C_{(i,j)}\}_{i=1j=1}^{M_{1}M_{2}}$, $0\leq C_{(i,j)}\leq L_{1}$.
5:for each $i=1,\cdots,M_{1}$ and $j=1,\cdots,M_{2}$ do
6:     if $C_{(i,j)}=0$ then
7:         $\overline{X}_{ij}=[\textbf{y}^{(i,j)}]E_{h_{i}}$
8:     else
9:         $\overline{X}_{ij}=[\textbf{y}^{(i,j)}]E_{0}$
10:     end if
11:     $C^{\prime}=H_{2}(d_{B},pp,\{\overline{X}_{ij}\}_{i=1j=1}^{M_{1}M_{2}})$.
12:     if $C=C^{\prime}$ then
13:         True.
14:     else
15:         False.
16:     end if
17:end for
18:True/False.
($\sigma$) $\leftarrow$ ${\sf Proxy~{}signature}$ $(m,SK_{B},d_{B},z_{B})$:- On input of a message $m\in\{0,1\}^{\star}$, $SK_{B}$, $d_{B}$ and $z_{B}$, B executes the following steps to generate a proxy signature $\sigma$ for $m$.
  1. 1.
    Computes $h^{\prime}={H_{1}}(z_{B},pp)$ and parses $h^{\prime}$ as integers $\{h_{i}^{\prime}\}_{i=1}^{M_{1}}$ for $0\leq h_{i}^{\prime}\leq L_{0}$.
  2. 2.
    For $i=1,\cdots,M_{1}$ and $j=1,\cdots,M_{2}$, chooses $\boldsymbol{\xi}^{(i,j)}\in_{R}[-(I_{0}+I_{1}),(I_{0}+I_{1})]^{n}$ and evaluates $Y_{(i,j)}=[\boldsymbol{\xi}^{(i,j)}]{\hat{E}_{h_{i}^{\prime}}}$.
  3. 3.
    Determines $D={H_{2}}(z_{B},pp,m,\{{Y}_{ij}\}_{i=1j=1}^{M_{1}M_{2}})$ and parses it to integers $\{D_{(i,j)}\}_{i=1j=1}^{M_{1}M_{2}}$ with $0\leq D_{(i,j)}\leq L_{1}$.
  4. 4.
    For $i=1,\cdots,M_{1}$ and $j=1,\cdots,M_{2}$ sets
    $\boldsymbol{\eta}^{(i,j)}=\begin{cases}\boldsymbol{\xi}^{(i,j)},&\text{if $D_{(i,j)}=0$}\\ \boldsymbol{\xi}^{(i,j)}+\textbf{b}^{(h^{\prime}_{i})},&\text{otherwise}\end{cases}$
  5. 5.
    If $\boldsymbol{\eta}^{(i,j)}\in[-I_{1},I_{1}]^{n}$ then performs the next step; otherwise, starts from step 2.
  6. 6.
    Outputs the signature as ${\sigma=\{D,{\{\boldsymbol{\eta}}^{(i,j)}\}_{i=1j=1}^{M_{1}M_{2}}}\}$.
Algorithm 6 Proxy signature
1:$(m,SK_{B},d_{B},z_{B})$, $m\in\{0,1\}^{\star}$
2:Computes $h^{\prime}={H_{1}}(z_{B},pp)$.
3:$h^{\prime}=\{h_{i}^{\prime}\}_{i=1}^{M_{1}}$, $h_{i}^{\prime}$’s are integers with $0\leq h_{i}^{\prime}\leq L_{0}$.
4:for each $i=1,\cdots,M_{1}$ and $j=1,\cdots,M_{2}$ do
5:     Chooses $\boldsymbol{\xi}^{(i,j)}\in_{R}[-(I_{0}+I_{1}),(I_{0}+I_{1})]^{n}$.
6:     Evaluates $Y_{(i,j)}=[\boldsymbol{\xi}^{(i,j)}]{\hat{E}_{h_{i}^{\prime}}}$.
7:end for
8:$D={H_{2}}(z_{B},pp,m,\{{Y}_{ij}\}_{i=1j=1}^{M_{1}M_{2}})$.
9:Parses $D$ to a integer set $\{D_{(i,j)}\}_{i=1j=1}^{M_{1}M_{2}}$ with $0\leq D_{(i,j)}\leq L_{1}$.
10:for each $i=1,2,\ldots,M_{1}$ and $j=1,2,\ldots,M_{2}$ do
11:     if $D_{(i,j)}=0$ then
12:         $\boldsymbol{\eta}^{(i,j)}=\boldsymbol{\xi}^{(i,j)}$
13:     else
14:         $\boldsymbol{\eta}^{(i,j)}=\boldsymbol{\xi}^{(i,j)}+\textbf{b}^{(h^{\prime}_{i})}$
15:     end if
16:     if $\boldsymbol{\eta}^{(i,j)}\in[-I_{1},I_{1}]^{n}$ then
17:         Perform the next step.
18:     else
19:         Start from Step 3.
20:     end if
21:end for
22:${\sigma=\{D,{\{\boldsymbol{\eta}}^{(i,j)}\}_{i=1j=1}^{M_{1}M_{2}}}\}$.
(0 or 1) $\leftarrow$ ${\sf Proxy~{}signature~{}verification}$ $(m,\sigma,PK_{A},PK_{B},d_{B},z_{B})$:- Using $(PK_{A},d_{B},z_{B})$, the verifier initially checks the validity of the warrant $(d_{B},z_{B})$ of B using the similar approach as discussed in proxy share verification. If the verification fails it outputs 00; otherwise, does the following:
  1. 1.
    Computes $h^{\prime}={H_{1}}(z_{B},pp)$, and parses $h^{\prime}$ and $D$ into integers as $\{h_{i}^{\prime}\}_{i=1}^{M_{1}}$ with $0\leq h_{i}^{\prime}\leq L_{0}$ and $\{D_{(i,j)}\}_{i=1j=1}^{M_{1}M_{2}}$ with $0\leq D_{(i,j)}\leq L_{1}$, respectively.
  2. 2.
    For $i=1,\cdots,M_{1}$ and $j=1,\cdots,M_{2}$ , sets
    $\overline{Y}_{(i,j)}=\begin{cases}[{\boldsymbol{\eta}}^{(i,j)}]{\hat{E}_{h_{i}^{\prime}}},&\text{if $D_{(i,j)}=0$}\\ [{\boldsymbol{\eta}}^{(i,j)}]{{E}_{0}},&\text{otherwise}\end{cases}$
  3. 3.
    Evaluates $D^{\prime}={H_{2}}(z_{B},pp,m,\{{\overline{Y}}_{ij}\}_{i=1j=1}^{M_{1}M_{2}})$.
  4. 4.
    If $D=D^{\prime}$, outputs 1; otherwise, outputs 0.
    Algorithm 7 Proxy signature verification
    1:$(m,\sigma,PK_{A},PK_{B},d_{B},z_{B})$.
    2:Evaluate $h=H_{1}(d_{B},pp)$.
    3:Parses $h$ as integers $\{h_{i}\}_{i=1}^{M_{1}}$, $0\leq h_{i}\leq L_{0}$.
    4:Parses $C$ as integers $\{C_{(i,j)}\}_{i=1j=1}^{M_{1}M_{2}}$, $0\leq C_{(i,j)}\leq L_{1}$.
    5:for each $i=1,\cdots,M_{1}$ and $j=1,\cdots,M_{2}$ do
    6:     if $C_{(i,j)}=0$ then
    7:         $\overline{X}_{ij}=[\textbf{y}^{(i,j)}]E_{h_{i}}$
    8:     else
    9:         $\overline{X}_{ij}=[\textbf{y}^{(i,j)}]E_{0}$
    10:     end if
    11:     $C^{\prime}=H_{2}(d_{B},pp,\{\overline{X}_{ij}\}_{i=1j=1}^{M_{1}M_{2}})$.
    12:     if $C=C^{\prime}$ then
    13:         Determines $h^{\prime}={H_{1}}(z_{B},pp)$.
    14:         Parses $h^{\prime}$ as $\{h_{i}^{\prime}\}_{i=1}^{M_{1}}$, where $h_{i}^{\prime}$’s are integers with $0\leq h_{i}^{\prime}\leq L_{0}$.
    15:         Parses $D$ as integers $\{D_{(i,j)}\}_{i=1j=1}^{M_{1}M_{2}}$ with $0\leq D_{(i,j)}\leq L_{1}$.
    16:         for each $i=1,\cdots,M_{1}$ and $j=1,\cdots,M_{2}$ do
    17:              if $D_{(i,j)}=0$ then
    18:                  $\overline{Y}_{(i,j)}=[{\boldsymbol{\eta}}^{(i,j)}]{\hat{E}_{h_{i}{\prime}}}$
    19:              else
    20:                  $\overline{Y}_{(i,j)}=[{\boldsymbol{\eta}}^{(i,j)}]{{E}_{0}}$
    21:              end if
    22:              $D^{\prime}={H_{2}}(z_{B},pp,m,\{{\overline{Y}}_{ij}\}_{i=1j=1}^{M_{1}M_{2}})$.
    23:              if $D=D^{\prime}$ then
    24:                  True.
    25:              else
    26:                  False.
    27:              end if
    28:         end for
    29:     else
    30:         False
    31:     end if
    32:end for
    33:True/False.
3.1 Correctness
To check the correctness of the proxy share and the proxy signature, we need to check the correctness of the equalities, $C=C^{\prime}$ and $D=D^{\prime}$, respectively.
$C=C^{\prime}$ will hold if and only if ${X_{ij}}=\overline{X}_{ij}$ for $1\leq i\leq M_{1}$, $1\leq j\leq M_{2}$.
Note that $X_{ij}=[\textbf{x}^{(i,j)}]E_{h_{i}}$ and
$\overline{X}_{ij}=\begin{cases}[\textbf{y}^{(i,j)}]E_{h_{i}},&\text{if $C_{(i,j)}=0$}\\ [\textbf{y}^{(i,j)}]E_{0},&\text{otherwise.}\end{cases}$
For $C_{(i,j)}=0$, $\textbf{y}^{(i,j)}=\textbf{x}^{(i,j)}$ which implies $\overline{X}_{ij}=[\textbf{y}^{(i,j)}]E_{h_{i}}=[\textbf{x}^{(i,j)}]E_{h_{i}}$. On the other hand, if $C_{(i,j)}\neq 0$ then $\textbf{y}^{(i,j)}=\textbf{x}^{(i,j)}+\textbf{a}^{(h_{i})}$, which implies $\overline{X}_{ij}=[\textbf{y}^{(i,j)}]E_{0}=[\textbf{x}^{(i,j)}+\textbf{a}^{(h_{i})}]E_{0}=[\textbf{x}^{(i,j)}]E_{h_{i}}$. Thus, the equality $C=C^{\prime}$ holds. Similarly, it can be proved that the equality $D=D^{\prime}$ holds.
4 Security analysis
In this section, we discuss the security of our proposed scheme.
  • Unforgeability of proxy signature: We now prove that our proposed isogeny based proxy signature scheme CSI-PS is uf-cma secure under the hardness of the group action inverse problem of isogeny. $H_{1}$ and $H_{2}$ are considered as random oracles. We show this by contradiction. If possible let there be an adversary $\mathcal{A}$ with a non-negligible advantage in the uf-cma game. Then we are going to show that using $\mathcal{A}$ an oracle machine $O^{\mathcal{A}}$ can be designed to break the isogeny problem by utilizing the outputs of $H_{1}$ and $H_{2}$. To do that, we present a series of games $Game^{0}$, $Game^{1}$, $Game^{2}$, $Game^{3}$, where $Game^{i}$ slightly modifies $Game^{i-1}$ for $i=1,2,3$. Let us denote the success probability of $\mathcal{A}$ in $Game^{i}$ as Prob[$Game^{i}$].
    1. 1.
      $\textbf{Game}^{0}$: $Game^{0}$ is exactly same as uf-cma game for proxy signature. Therefore, $Adv_{\mathcal{A}}^{Exp^{uf-cma}_{proxy(1^{\lambda})}}$ = Prob[$Exp^{uf-cma}_{proxy(1^{\lambda})}$ = 1] = Prob[$Game^{0}$].
    2. 2.
      $\textbf{Game}^{1}$: It is similar to $Game^{0}$, except that $O^{\mathcal{A}}$ replaces output of the the hash function $H_{1}$ on $(z_{B},pp)$ by a random string of length $M_{1}$ from $(\mathbb{Z}_{{L_{0}}+1})^{M_{1}}$. If $|Prob(Game^{1})-Prob(Game^{0})|$ is non-negligible then $\mathcal{A}$ will be able to distinguish the output distribution of $H_{1}$ which is impossible since $H_{1}$ is chosen as random oracle. Hence $|Prob(Game^{1})-Prob(Game^{0})|$ should be negligible, say bounded by $\epsilon_{1}(\lambda)$.
    3. 3.
      $\textbf{Game}^{2}$: $Game^{2}$ is similar to $Game^{1}$, except that $O^{\mathcal{A}}$ substitutes output of the hash function $H_{2}$ by a random string chosen from $(\mathbb{Z}_{{L_{1}}+1})^{M_{1}M_{2}}$. By a similar argument as discussed in $Game^{1}$, $|Prob(Game^{2})-Prob(Game^{1})|$leq $\epsilon_{2}(\lambda)$ (negligible).
    4. 4.
      $\textbf{Game}^{3}$: This is similar to $Game^{2}$, except that during the challenge phase when $\mathcal{A}$ tries to forge a signature of a message $m^{*}$, $O^{\mathcal{A}}$ substitutes $H_{1}$ output by a random tuple from $(\mathbb{Z}_{{L_{0}}+1})^{M_{1}}$ and substitute $H_{2}$ by a random tuple from $(\mathbb{Z}_{{L_{1}}+1})^{M_{1}M_{2}}$. Using the similar argument as discussed in $Game^{1}$, we may conclude that $|Prob(Game^{3})-Prob(Game^{2})|$leq $\epsilon_{3}(\lambda)$ (negligible).
    Now $|Prob(Game^{3})-Adv_{\mathcal{A}}^{Exp^{uf-cma}_{\textit{CSI-PS}(1^{\lambda})}}|$= $|Prob(Game^{3})-Prob(Game^{0})|$
    $\leq$ $|Prob(Game^{3})-Prob(Game^{2})|$+$|Prob(Game^{2})-Prob(Game^{1})|$+
    $|Prob(Game^{1})-Prob(Game^{0})|$leq $\epsilon_{1}(\lambda)$+$\epsilon_{2}(\lambda)$+$\epsilon_{3}(\lambda)$=$\epsilon(\lambda)$ (say).
    Therefore, the success probability of $\mathcal{A}$ in $Game^{3}$ is close to $Adv_{AD}^{Exp^{uf-cma}_{\textit{CSI-PS}(1^{\lambda})}}$. Since we have assumed that $Adv_{\mathcal{A}}^{Exp^{uf-cma}_{\textit{CSI-PS}(1^{\lambda})}}$ is non-negligible we can ensure that $Prob(Game^{3})$ is also non-negligible. Thereby, with the held of $\mathcal{A}$ and controlling $H_{1}$, $H_{2}$, the oracle machine $O^{\mathcal{A}}$ can generate two valid transcripts ($(\{{Y}_{ij}\}$, $D^{(1)}$, ${\boldsymbol{\eta}_{1}^{(i,j)}})$ and ($(\{{Y}_{ij}\},D^{(2)},{\boldsymbol{\eta}_{2}^{(i,j)}})$ where $D^{(1)}_{(i,j)}=0$ and $D^{(2)}_{(i,j)}\neq 0$ $\forall$ $i$, $j$. Note that $\boldsymbol{\eta}_{1}^{(i,j)}=\boldsymbol{\xi}^{(i,j)}$ since $D^{(1)}_{(i,j)}=0$ and $\boldsymbol{\eta}_{2}^{(i,j)}=\boldsymbol{\xi}^{(i,j)}+\textbf{b}^{\left(h^{\prime}_{i}\right)}$ since $D^{(2)}_{(i,j)}\neq 0$. Utilizing $\boldsymbol{\eta}_{1}^{(i,j)}$ and $\boldsymbol{\eta}_{2}^{(i,j)}$, $O^{\mathcal{A}}$ can easily extract the secret $SK_{B}=\{B^{i}\}_{i=1}^{L_{0}}$. Therefore, the group action inverse problem ($\sf{GAIP}$) between two elliptic curves is broken which leads to a contradiction. Thus, the advantage of $\mathcal{A}$ in the uf-cma game can not be non-negligible. In other words, our scheme CSI-PS is uf-cma secure.
    We now discuss the other security properties of our proposed design CSI-PS.
  • Identifiability: Given a signature $\sigma$ with the proxy share $z_{B}$, the verifier finds $(d_{B},z_{B})$ from the bulletin board and verifies the correctness of $z_{B}$. If it is valid, then from $d_{B}$, the identity of the proxy signer is detected.
  • Undeniability: At the stage of the verification process of the proxy signature, the proxy signer cannot deny their generation of the proxy signature. Here, the proxy signature contains ${\{\mathbf{\eta}}^{(i,j)}\}_{i=1j=1}^{M_{1}L_{1}}$, which is constructed by using secret-key of the proxy signer. Hence, after generating the proxy signature, the proxy signer can not disagree with their signature generation.
  • Verifiability: In the verification process, the verifier has to be convinced about the agreement of the original signature and the proxy signer from the proxy signature. To generate the warrant $z_{B}$, the secret key of the original signer has been used. Thus, by verifying the correctness of $z_{B}$ using $PK_{A}$, the verifier is convinced about the agreement.
  • Distinguishability: Anyone authorized to verify a signature can distinguish between a proxy signature and a normal signature. In the isogeny-based signature scheme Seasign [26], the signer generates a signature on a message using their own secret key, and the verifier confirms the signature using the signer’s public key. On the other hand, in a proxy signature scheme, the proxy signer uses their own proxy share $z_{B}$ (generated by the original signer) and own secret key. The verifier then checks the signature using the proxy share and the public information of both the original signer and the proxy signer. Therefore, the proxy signature and the ordinary signature are distinguishable due to their different structures.
  • Secrecy: Given the proxy share $z_{B}$, it is impossible to derive the original signer’s secret key. Deriving the secret key from the warrant would require solving the GAIP problem (see Definition 2) for the instances $(X_{ij},E_{h_{i}})$, where $i=1,2,\ldots,M_{1}$ and $j=1,2,\ldots,M_{2}$. Thus, extracting the original signer’s secret key is computationally hard. Similarly, it is difficult to extract the proxy signer’s secret key from a proxy signature.
  • Prevention of misuse: The proxy signer can sign a message using the warrant $(d_{B},z_{B})$ assigned specifically to them. This warrant includes their identity and other necessary information. Therefore, any misuse of the proxy signature can be traced back to the proxy signer through the warrant.
  • Revocability: Once the time period involved in the warrant $(d_{B},z_{B})$ is over, automatically the signing authority of the proxy signer is removed. Besides, the original signer can broadcast the announcement of invalidation of the warrant $(d_{B},z_{B})$ by putting a signed message on the bulletin board.
5 Complexity analysis of CSI-PS
We describe below the communication and computation complexity of our proposed CSI-PS. The public key sizes of the original signer and the proxy signer, the proxy share size, and the proxy signature size all contribute to the scheme’s communication costs. Each signer’s public key is made up of $L_{0}$ number of elliptic curves. According to Theorem 1, any elliptic curve can be represented by a single coefficient in $\mathbb{F}_{p}$ after class group action. Additionally, proxy share generation and proxy signature generation both contains one hash output and $M_{1}M_{2}$ vectors from $[-I_{1},I_{1}]^{n}$. In our scheme, we have two hash functions $H_{1}$ and $H_{2}$. In proxy share generation and proxy signature generation algorithms four hash outputs $h,h^{\prime}$, $C$ and $D$ have been used. The lengths of $h$ and $h^{\prime}$ are $M_{1}\left\lceil{log(L_{0}+1)}\right\rceil$ bits. The lengths of $C$ and $D$ are $M_{1}M_{2}\left\lceil{log(L_{1}+1)}\right\rceil$ bits. The communication complexity of our proposed scheme CSI-PS is given in Table 1.
To achieve $\lambda$ bits of security we may take $M_{1}\left\lceil{log(L_{0}+1)}\right\rceil\geq\lambda$ and $M_{1}M_{2}\left\lceil{log(L_{1}+1)}\right\rceil\geq\lambda$. Since $L_{0}=2^{\gamma_{0}}-1$ and $L_{1}=2^{\gamma_{1}}-1$, we conclude that $M_{1}{\gamma_{0}}\geq\lambda$ and $M_{1}M_{2}{\gamma_{1}}\geq\lambda$. In particular, we select our own parameters $M_{1}$ and $M_{2}$ as $M_{1}\gamma_{0}=\lambda$, $M_{1}M_{2}\gamma_{1}=\lambda$. Hence, $M_{1}=\frac{\lambda}{\gamma_{0}}$, $M_{1}M_{2}=\frac{\lambda}{\gamma_{1}}$ and $M_{2}=\frac{\gamma_{0}}{\gamma_{1}}$. We choose $\alpha_{0}=nM_{1}L_{1}$. Castryck et al. [23] proposed parameters for CSIDH-512 to achieve AES-128 bit security. CSIDH-512 includes the prime of the form $p=4l_{1}l_{2}\ldots l_{n}-1$ where, $n=74$, $I_{73}=373$ and $I_{74}=587$, the key bound $m$ is taken as 5 (See subsection 2.3 for details of the parameters). In our scheme $n=74$, $I_{0}=5$. In order to calculate the sizes of the keys, proxy share and proxy signature, we rely on the different choices of parameters $(\gamma_{0},\gamma_{1}M_{1},M_{2},L_{0},L_{1})$ as proposed in [29] for $\lambda=128$-bit security level. For some choices of parameters the results are documented in Table 2.
Table 1: The sizes of public keys, secret keys, proxy share, and proxy signature.
Data Size (bits)
$PK_{A}$ $L_{0}\left\lceil{log(p)}\right\rceil$
$SK_{A}$ $L_{0}\left\lceil{nlog(2I_{0}+1)}\right\rceil$
$PK_{B}$ $L_{0}\left\lceil{log(p)}\right\rceil$
$SK_{B}$ $L_{0}\left\lceil{nlog(2I_{0}+1)}\right\rceil$
Proxy share($z_{B}$) $M_{1}M_{2}\left\lceil{log(L_{1}+1)}\right\rceil$ + $M_{1}M_{2}\left\lceil{nlog(2I_{1}+1)}\right\rceil$
Proxy Signature($\sigma$) $M_{1}M_{2}\left\lceil{log(L_{1}+1)}\right\rceil$ + $M_{1}M_{2}\left\lceil{nlog(2I_{1}+1)}\right\rceil$
Table 2: Communication cost and secret key size for different parameters for 128 bit security parameters
$(\gamma_{0},\gamma_{1},M_{1},M_{2},L_{0},L_{1})$ Public key size Secret key size Proxy share or signature size
(1, 1, 128, 1, 1, 1) 63.75B 32B 19584B
(1, 8, 16, 1, 1, 255) 63.75B 32B 3198B
(8, 16, 16, 1, 255, 65535) 16256.25B 8160B 4402B
(16, 8, 8, 2, 65535, 255) 4177856.25B 2097120B 3052B
6 Concrete instantiation
In this section, we discuss and analyze a concrete instantiation of CSI-PS for $128$-bit security level with parameter set $(16,8,8,2,65535,255)$ Suppose a government officer A (original signer) is the only person authorized to sign a specific document. In case $A$ is on the leave, they must appoint a representative officer B (proxy signer) to delegate the signing authority. We demonstrate how B will obtain signing rights from A using a toy example.
  • Here $I_{0}=5$, $n=74$, $\gamma_{0}=16$, $\gamma_{1}=8$, $M_{1}=8$, $M_{2}=2$, $L_{0}=2^{16}-1=65535$, $L_{1}=2^{8}-1=255$, $I_{1}=754800$
  • A selects the vectors $\textbf{a}^{1},\cdots,\textbf{a}^{65535}$ randomly from $[-5,5]^{74}$.
  • A computes $E_{1},\cdots,E_{65535}$, where $E_{i}=[\textbf{a}^{(i)}]E_{0}$, $i=1,\cdots,65535$.
  • B selects the vectors $\textbf{b}^{1},\cdots,\textbf{b}^{65535}$ randomly from $[-5,5]^{74}$.
  • B computes $\hat{E_{1}},\cdots,\hat{E}_{65535}$, where $\hat{E_{i}}=[\textbf{b}^{(i)}]E_{0}$, $i=1,\cdots,65535$.
$SK_{A}=\textbf{a}^{1},\cdots,\textbf{a}^{65535}$, $PK_{A}=E_{1},\cdots,E_{65535}$ and $SK_{B}=\textbf{b}^{1},\cdots,\textbf{b}^{65535}$, $PK_{B}=\hat{E_{1}},\cdots,\hat{E}_{65535}$. The original signer will now perform the following actions.
  • $h=H_{1}(d_{B},pp)={h_{1},\cdots,h_{8}}$, where $h_{1},\cdots,h_{8}\in\mathbb{Z}_{65536}$.
  • Randomly selects, $\textbf{x}^{(1,1)},\textbf{x}^{(1,2)},\textbf{x}^{(2,1)},\textbf{x}^{(2,2)},\cdots,\textbf{x}^{(8,1)},\textbf{x}^{(8,2)}$ from $[754795,754805]^{74}$
  • Evaluates $X_{11}=[\textbf{x}^{(1,1)}]E_{h_{1}}$, $X_{12}=[\textbf{x}^{(1,2)}]E_{h_{1}}$, $X_{21}=[\textbf{x}^{(2,1)}]E_{h_{2}}$, $X_{22}=[\textbf{x}^{(2,2)}]E_{h_{2}}$, $\cdots$, $X_{81}=[\textbf{x}^{(8,1)}]E_{h_{8}}$, $X_{ij}=[\textbf{x}^{(8,2)}]E_{h_{8}}$.
  • $C=H_{2}(d_{B},pp,\{{X}_{ij}\}_{i=1j=1}^{8~{}~{}2})$= $\{C_{(1,1)},C_{(1,2)},C_{(2,1)},C_{(2,2)},\cdots,C_{(8,1)},C_{(8,2)}\}\subset\mathbb{Z}_{256}$.
  • Computes $\textbf{y}^{(1,1)},\textbf{y}^{(1,2)},\textbf{y}^{(2,1)},\textbf{y}^{(2,2)},\cdots,\textbf{y}^{(8,1)},\textbf{y}^{(8,2)}$ according to Algorithm 4.
Therefore the proxy share is $z_{B}=(C,\{\mathbf{y}^{(i,j)}\}^{8~{}~{}2}_{i=1j=1})$. Using Algorithm 5, the representative proxy signer can validate the proxy share using its own credentials. In the following, signing rights of the original signer might be assigned to the representative signer B. B can now use the signing rights to sign the respective documents on the behalf of the original signer A. We now demonstrate how the representative officer can sign a document.
  • Evaluates $h^{\prime}={H_{1}}(z_{B},pp)=\{h_{1}^{\prime},\cdots,h_{8}^{\prime}\}$, where $h_{1}^{\prime},\cdots,h_{8}^{\prime}\in\mathbb{Z}_{65536}$.
  • Randomly selects, $\boldsymbol{\xi}^{(1,1)},\boldsymbol{\xi}^{(1,2)},\boldsymbol{\xi}^{(2,1)},\boldsymbol{\xi}^{(2,2)},\cdots,\boldsymbol{\xi}^{(8,1)},\boldsymbol{\xi}^{(8,2)}$ from $[754795,754805]^{74}$.
  • Computes the following elliptic curves, $Y_{11}=[\textbf{x}^{(1,1)}]\hat{E}_{h_{1}^{\prime}}$, $Y_{12}=[\boldsymbol{\xi}^{(1,2)}]\hat{E}_{h_{1}^{\prime}}$, $Y_{21}=[\boldsymbol{\xi}^{(2,1)}]\hat{E}_{h_{2}^{\prime}}$, $Y_{22}=[\boldsymbol{\xi}^{(2,2)}]\hat{E}_{h_{2}^{\prime}}$, $\cdots$, $Y_{81}=[\boldsymbol{\xi}^{(8,1)}]\hat{E}_{h_{8}^{\prime}}$, $Y_{ij}=[\boldsymbol{\xi}^{(8,2)}]\hat{E}_{h_{8}^{\prime}}$.
  • Assume that $m$ represents the content of the document that the representative officer wishes to sign. To proceed further he calculates the hash value, $D=H_{2}(z_{B},pp,m,\{{Y}_{ij}\}_{i=1j=1}^{8~{}~{}2})$= $\{D_{(1,1)},D_{(1,2)},D_{(2,1)},D_{(2,2)},\cdots,D_{(8,1)},D_{(8,2)}\}\subset\mathbb{Z}_{256}$.
  • Computes $\boldsymbol{\eta}^{(1,1)},\boldsymbol{\eta}^{(1,2)},\boldsymbol{\eta}^{(2,1)},\boldsymbol{\eta}^{(2,2)},\cdots,\boldsymbol{\eta}^{(8,1)},\boldsymbol{\eta}^{(8,2)}$ by the method given in Algorithm 6 and checks whether $\boldsymbol{\eta}^{(1,1)},\boldsymbol{\eta}^{(1,2)},\boldsymbol{\eta}^{(2,1)},\boldsymbol{\eta}^{(2,2)},\cdots,\boldsymbol{\eta}^{(8,1)},\boldsymbol{\eta}^{(8,2)}\in[-754800,754800]^{74}$.
Finally the proxy signer outputs the proxy signature ${\sigma=\{D,{\{\boldsymbol{\eta}}^{(i,j)}\}_{i=1j=1}^{8~{}~{}2}}\}$. The designated verifier can verify the proxy signature using the Algorithm 7.
We now discuss the computation complexity of CSI-PS for $128$-bit security level corresponding to aforementioned parameter set. The most computational intensive operation in CSI-PS is the class group computation. In the first step, we estimated the cost of one class group computation in terms of field multiplications. In the following, we compute the computational complexity of public key generation, proxy share generation, and proxy signature generation in terms of total number of field multiplications required. We also calculated the time complexity of CSI-PS by estimating the run time and cpu clock cycles for public key generation, proxy share generation, and proxy signature generation. We ran the C script [^1] https://yx7.cc/code/csidh/csidh-latest.tar.xz on a workstation having Intel Core i5-8700 2.40GHz processor with 8GB of RAM and Linux Lite v5.2 OS. The results of our experiments are summarized in Table 3.
Table 3: Computation complexity and time complexity of CSI-PS for 128 bit security level
Clock cycle Time (in seconds) Field multiplications
CSIDH-512 (one class group computation) $45.7\times 10^{6}$ $0.021624$ $282,424$
Public Key $29.94\times 10^{11}$ $1422.12$ $18,508,656,840$
Proxy Share $73.12\times 10^{8}$ $0.3472$ $4,518,784$
Proxy Signature $73.12\times 10^{8}$ $0.3472$ $4,518,784$
7 Conclusion
This paper presents a promising approach to delegate the signing rights of an original signer to proxy signers using isogeny-based cryptography. The isogeny-based signature scheme SeaSign [26] serves as the fundamental building block for our scheme. Our proposed scheme achieves uf-cma security under the assumption that the group action inverse problem ($\sf{GAIP}$) between two elliptic curves is hard. Additionally, our scheme satisfies other security properties of a proxy signature. In our design, the proxy signature is $M_{1}M_{2}\left\lceil{\log(L_{1}+1)}\right\rceil$ + $M_{1}M_{2}\left\lceil{n\log(2I_{1}+1)}\right\rceil$ bits in size. Through calculations, we conclude that the proxy signature size is 3052B for a 128-bit security level with a certain parameter choice. Constructing the proxy signature requires 16 class group operations or 4,518,784 field multiplications. In particular, our proposed proxy signature scheme, CSI-PS, is the first of its kind using isogeny-based cryptography.
Declarations
Funding Details
This work was supported by the “International Mathematical Union (IMU) and the Graduate Assistantships in Developing Countries (GRAID) Program”.
Conflict of interest
The authors state that they have not known competing financial interests or personal connections that may seem to have influenced the work described in this study.
Data availability
Data sharing is not applicable to this article as no new data were generated or analyzed to support this research.